OpenStudy (anonymous):

Factorize : 4a2-9b2-2a-3b.

OpenStudy (lgbasallote):

groupp them first \[\implies (4a^2 - 9b^2) - (2a + 3b)\] now factor out \[\implies (2a+3b)(2a-3b) - (2a+3b)\] does that help @basith ?

OpenStudy (anonymous):

after this ???? @lgbasallote

OpenStudy (ajprincess):

Nw factor out (2a+3b) nd see vat u get.

OpenStudy (anonymous):

hw @ajprincess

OpenStudy (ajprincess):

\((2a+3b)(2a-3b)-(2a+3b)\) assume (2a+3b)=x. and (2a-3b)=y then u hav xy-x Vat do u get when u factor out x?

OpenStudy (anonymous):

how to factor....i dont remember hw to ...pls...show....

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

xy-x If you factor out x u get x(y-1). K? nw can u find ur answer?

OpenStudy (anonymous):

cn u show pls......i get a little confused when doing the problem..@ajprincess

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

Sorry hav to go nw. Wil come back to u in few mnts.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@Neemo cn u pls..HELP

OpenStudy (ajprincess):

xy-x xy means x*y and x means x*1 So when you take x from xy only y remains and when you take x from x vat remains is 1. is that clear?

OpenStudy (anonymous):

i dont hv a doubt about tht.....but cn u do the problem in equation form.....

OpenStudy (anonymous):

pls...write it down in the equation form

OpenStudy (ajprincess):

x^2-5x+6=0 Something like this?

OpenStudy (ajprincess):

@basith

OpenStudy (anonymous):

cn u write for me the full answer like...this \[23 \sqrt{2}\]

OpenStudy (anonymous):

i meant the writings shud be like that...cn u????

OpenStudy (ajprincess):

u want me to write in latex?

OpenStudy (ajprincess):

\(xy-x\) when u factor out x u get \(x(y-1)\). that is \(xy-x=x(y-1)\) K?

OpenStudy (anonymous):

ni in X and Y ..........can u show me the answer of the question...imean by sustituting X and Y

OpenStudy (ajprincess):

Substitute \((2a+3b)\) for \(x\) and \((2a-3b)\) for y in \(x(y-1)\) vat do u get when u substitute?

OpenStudy (ajprincess):

@basith

OpenStudy (anonymous):

\[2a+3b (?-1)\]

OpenStudy (anonymous):

\[2a+3b(2a-3b-1) ..........like this????\]

OpenStudy (ajprincess):

ya.

OpenStudy (ajprincess):

Is it clear nw?

OpenStudy (anonymous):

then wat :( ????